# Tac2Motion: Contact-Aware Reinforcement Learning with Tactile Feedback for Robotic Hand Manipulation

### v1

@@ I Introduction
Robotic hand manipulation often requires contact-rich dexterity, which has long remained an exceptional challenge due to complex high dimensionality, lack of sensing modality, and limited generalization. Rapid progress in Reinforcement Learning (RL) has accelerated a major step to overcome the above issues, which in turn has enhanced robotic in-hand manipulation capabilities [1]. In addition, visual feedback often assists in-hand manipulation control policies, showing outstanding performances [2, 3].
Even though using visual sensory feedback into RL is widely used, it is still limited by occlusion from fingers. To overcome this, one promising approach is to add extra cameras [4], which is unrealistic in practice. Unlike the vision system, tactile sensory feedback is free from the occlusion issue and is a more natural way to obtain useful contact information efficiently. Inspired by this, human demonstrations with the tactile sensor technology (e.g. GelSight [5]) have enabled learning human’s dexterity efficiently [6, 7]. However, employing human data is quite expensive due to the kinematic and dynamic embedding problem [8].
To get around these barriers, model-free RL has provided a viable alternative for learning contact-rich dexterity, showing meaningful performance in multi-finger manipulation domain with tactile sensing [9, 10, 11, 12, 13, 14]. Despite the remarkable progress in RL-based dexterous frameworks with tactile sensing, many of them have often neglected the benefits from using tactile sensors in shaping reward functions for learning a dexterous control policy, and the related research has not yet been fully investigated.
@@ II-A Learning Control Policy
During training the control policy, agents observe $\bm{o}_{t}$ including joint position, $\bm{q}_{t}$, joint angular velocity, $\dot{\bm{q}}_{t}$, previous joint target position, $\bm{q}^{d}_{t-1}$, tactile embedding, $\bm{z}_{t}$, and the center position of a lid, $\bm{p}$ at each time step, $t$ as shown in Fig. 1. Action space includes a relative desired joint position, $\bm{a}_{t}$, and the final target joint position is decided by adding $\bm{a}_{t}$} to $\bm{q}^{d}_{t-1}$. To avoid jerky motions, we use Exponential Moving Average (EMA) function to make the target values smooth, $\bm{q}^{d}_{t}=\bm{q}^{d}_{t-1}+\eta\textnormal{EMA}(\bm{a}_{t})$ where $\eta$ is a scaling parameter [15]. Lastly, we utilize PPO to train the control policy, accompanying asymmetric learning and actor-critic structure.
@@ II-B Reward Design
3) Rotation Reward, (RR): We newly define the torsional reward designed to promote grasping and rotation at the same time, extending the previous twisting reward proposed [15]:
@@ II-C Penalty Design
Adding penalty terms helps to discourage the undesired behaviors such as incorrect rotation direction and jerky finger motions. We also consider the auxiliary penalty terms used in [15] in the following:
@@ III Simulation
In this section, we carry out several simulations to investigate the performance of the learned contact-aware in-hand manipulation in the lid-removal scenario. We evaluate the proposed method based on the following metrics: Rotation Score (RS), Rotation Time (RT), and Success Rate (SR). We then compare the proposed method with the baseline provided by [15] and conduct an ablation study to evaluate the combination of reward functions.
@@ III-A Simulation Setups
We conduct the simulation in IsaacGym by modifying the code in [15]. We use multi-fingered robotic hand, Shadow Robot to assess the performance of each method and also leverage the guided contact fixtures suggested in [15] on the lids as shown in Fig. A.1. Different geometries (e.g. cylinder, square, and hexagon), sizes, frictions of lids are considered in our simulations. We train each method with $10240$ environments and batch sizes, $10240$ with one single NVIDIA GeForce RTX 4090 GPU, which takes $20$ minutes for the entire training. Afterwards, we test the learned models on $100$ agents to show general performances. Note that the tactile embedding is frozen on training and testing.
@@ III-B Virtual Torque
Simulating patch contact interactions and their effects (e.g. increase of contact friction) is non-trivial and highly computational. To resolve this, we introduce a virtual torque on a lid that corresponds to the variation of the contact friction caused by the patch contact interaction. The higher the normal magnitude of each tactile sensor, the more it is likely to increase the friction, which in turn generates a stronger torque applied to the lid. This technique is distinguished from the penetration-based friction calculation that is commonly used in the simulation, enabling the practical implementation of soft finger model [16].
@@ III-D Generalization of Tac2Motion
We compare Tac2Motion with several methods to verify its efficacy, including baseline [15], a couple of combinations of reward functions. We simulate each method on three types of lids, such as cylinder, square, and hexagon, to observe the generalization performance. The multiple metrics to evaluate the performance are leveraged: Rotation Score (RS), Rotation Time (RT), and Success Rate (SR). RS is to analyze the angle changes of the lid per each step, and RT is the average time for one revolution of the lid. We also define success and failure based on whether the lid completes one full revolution within certain limits, $2.5$ [s], $5.0$ [s], and $3.5$ [s] corresponding to cylinder, square, and hexagon, respectively. As summarized in Table. I, Tac2Motion shows the superior performances across all metrics. RS and RT indicate that Tac2Motion achieves the fastest rotation of the lid, and SR shows the most robust rotational performance compared to the baseline. These advantages arise because Tac2Motion encourages the agent to behave with firm grasping and smooth finger gaiting simultaneously during rotation, while the baseline primarily focuses on rotation performance. In an ablation study, it is observed that applying CPR reward (CPR+RR) is highly beneficial for removing the lid since maintaining grasping of the lid during rotating ensures stable finger behavior. In contrast, using only CRR (CRR+RR) does not lead to meaningful final performance according to the success rate and rotation score as presented in Table. I. Therefore, the results of the simulation support that contact-aware reward design improves the performance of contact-rich in-hand manipulation.

### v2

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Robotic hand manipulation often requires contact-rich dexterity, which has long remained an exceptional challenge due to complex high dimensionality, lack of sensing modality, and limited generalization. Rapid progress in Reinforcement Learning (RL) has accelerated a major step to overcome the above issues, which in turn has enhanced robotic in-hand manipulation capabilities [yin2025learninginhandtranslationusing]. In addition, visual feedback often assists in-hand manipulation control policies, showing outstanding performances [SridharICRA2023, HoloSridharICRA2023].
Even though using visual sensory feedback into RL is widely used, it is still limited by occlusion from fingers. To overcome this, one promising approach is to add extra cameras [li2025dropdexterousreorientationonline], which is unrealistic in practice. Unlike the vision system, tactile sensory feedback is free from the occlusion issue and is a more natural way to obtain useful contact information efficiently. Inspired by this, human demonstrations with the tactile sensor technology (e.g. GelSight [YuanGelsight2017]) have enabled learning human’s dexterity efficiently [iyer2024open, HuangTMech2025]. However, employing human data is quite expensive due to the kinematic and dynamic embedding problem [pmlr-v164-chen22a].
To get around these barriers, model-free RL has provided a viable alternative for learning contact-rich dexterity, showing meaningful performance in multi-finger manipulation domain with tactile sensing [ChenTransactions2024, OpenAI2019, openai2019rubiks, ZhuICRA2019, YangRAL2023, GuzeyICRA2024]. Despite the remarkable progress in RL-based dexterous frameworks with tactile sensing, many of them have often neglected the benefits from using tactile sensors in shaping reward functions for learning a dexterous control policy, and the related research has not yet been fully investigated.
@@ II-A Learning Control Policy
During training the control policy, agents observe $\bm{o}_{t}$ including joint position, $\bm{q}_{t}$, joint angular velocity, $\dot{\bm{q}}_{t}$, previous joint target position, $\bm{q}^{d}_{t-1}$, tactile embedding, $\bm{z}_{t}$, and the center position of a lid, $\bm{p}$ at each time step, $t$ as shown in Fig. 1. Action space includes a relative desired joint position, $\bm{a}_{t}$, and the final target joint position is decided by adding $\bm{a}_{t}$} to $\bm{q}^{d}_{t-1}$. To avoid jerky motions, we use Exponential Moving Average (EMA) function to make the target values smooth, $\bm{q}^{d}_{t}=\bm{q}^{d}_{t-1}+\eta\textnormal{EMA}(\bm{a}_{t})$ where $\eta$ is a scaling parameter [pmlr-v270-lin25c]. Lastly, we utilize PPO to train the control policy, accompanying asymmetric learning and actor-critic structure.
@@ II-B Reward Design
3) Rotation Reward, (RR): We newly define the torsional reward designed to promote grasping and rotation at the same time, extending the previous twisting reward proposed [pmlr-v270-lin25c]:
@@ II-C Penalty Design
Adding penalty terms helps to discourage the undesired behaviors such as incorrect rotation direction and jerky finger motions. We also consider the auxiliary penalty terms used in [pmlr-v270-lin25c] in the following:
@@ III Simulation
In this section, we carry out several simulations to investigate the performance of the learned contact-aware in-hand manipulation in the lid-removal scenario. We evaluate the proposed method based on the following metrics: Rotation Score (RS), Rotation Time (RT), and Success Rate (SR). We then compare the proposed method with the baseline provided by [pmlr-v270-lin25c] and conduct an ablation study to evaluate the combination of reward functions.
@@ III-A Simulation Setups
We conduct the simulation in IsaacGym by modifying the code in [pmlr-v270-lin25c]. We use multi-fingered robotic hand, Shadow Robot to assess the performance of each method and also leverage the guided contact fixtures suggested in [pmlr-v270-lin25c] on the lids as shown in Fig. A.1. Different geometries (e.g. cylinder, square, and hexagon), sizes, frictions of lids are considered in our simulations. We train each method with $10240$ environments and batch sizes, $10240$ with one single NVIDIA GeForce RTX 4090 GPU, which takes $20$ minutes for the entire training. Afterwards, we test the learned models on $100$ agents to show general performances. Note that the tactile embedding is frozen on training and testing.
@@ III-B Virtual Torque
Simulating patch contact interactions and their effects (e.g. increase of contact friction) is non-trivial and highly computational. To resolve this, we introduce a virtual torque on a lid that corresponds to the variation of the contact friction caused by the patch contact interaction. The higher the normal magnitude of each tactile sensor, the more it is likely to increase the friction, which in turn generates a stronger torque applied to the lid. This technique is distinguished from the penetration-based friction calculation that is commonly used in the simulation, enabling the practical implementation of soft finger model [Prattichizzo2016].
@@ III-D Generalization of Tac2Motion
We compare Tac2Motion with several methods to verify its efficacy, including baseline [pmlr-v270-lin25c], a couple of combinations of reward functions. We simulate each method on three types of lids, such as cylinder, square, and hexagon, to observe the generalization performance. The multiple metrics to evaluate the performance are leveraged: Rotation Score (RS), Rotation Time (RT), and Success Rate (SR). RS is to analyze the angle changes of the lid per each step, and RT is the average time for one revolution of the lid. We also define success and failure based on whether the lid completes one full revolution within certain limits, $2.5$ [s], $5.0$ [s], and $3.5$ [s] corresponding to cylinder, square, and hexagon, respectively. As summarized in Table. I, Tac2Motion shows the superior performances across all metrics. RS and RT indicate that Tac2Motion achieves the fastest rotation of the lid, and SR shows the most robust rotational performance compared to the baseline. These advantages arise because Tac2Motion encourages the agent to behave with firm grasping and smooth finger gaiting simultaneously during rotation, while the baseline primarily focuses on rotation performance. In an ablation study, it is observed that applying CPR reward (CPR+RR) is highly beneficial for removing the lid since maintaining grasping of the lid during rotating ensures stable finger behavior. In contrast, using only CRR (CRR+RR) does not lead to meaningful final performance according to the success rate and rotation score as presented in Table. I. Therefore, the results of the simulation support that contact-aware reward design improves the performance of contact-rich in-hand manipulation.